Dad is 6 times as old as Michael and he will be 4 times as old as Michael in 4 years time. If dad is less than 60 years of age, how old will he be in 6 year's time?
Answer by ankor@dixie-net.com(22740)

d = Dad's age
m - Michael's age
:
Dad is 6 times as old as Michael
d = 6m
:
and he will be 4 times as old as Michael in 4 years time.
d + 4 = 4(m+4)
d + 4 = 4m + 16
d = 4m + 16 - 4
d = 4m + 12
replace d with 6m
6m = 4m + 12
6m - 4m = 12
2m = 12
m = 6 yrs is M's age
then
6*6 = 36 yrs is Dad's age
:
If dad is less than 60 years of age, how old will he be in 6 year's time?
36 + 6 = 42 yrs